OpenStudy (meshaw24):

Which two equations are equivalent? A y = (x + 4)2 y = x2 + 16 B y = (x – 3)2 y = x2 – 6x + 9 C y = (x + 2)2 y = x2 + 2x + 2 D y = (x – 6)2 y = x2 – 12x – 36

OpenStudy (mww):

expand the perfect squares and see what you get. Remember that \[(x+y)^2 = x^2 +2xy + y^2\] and \[(x-y)^2 = x^2 -2xy +y^2\] If the expanded result is the same as the one below, they are equivalent

OpenStudy (peachpi):

multiply the top equation and see if you get the second. For example in A, is (x+4)(x + 4) the same as x^2 + 16?